# IMPROVE: Iterative Model Pipeline Refinement and Optimization Leveraging LLM Agents

### v1

@@ 3.3 IMPROVE
To evaluate the Iterative Refinement strategy, we developed IMPROVE, a multi-agent end-to-end LLM framework that autonomously generates and optimizes an object classification pipeline. IMPROVE requires only a dataset as input and independently manages the entire process, from pipeline generation to model training, eliminating the need for any human intervention. We provide a diagram of the workflow of IMPROVE in Figure 1.
@@
IMPROVE consists of the following key agents:
  • Project Architect: This agent is responsible for the initial analysis of the dataset, evaluating its size, structure, and any supplementary information provided. Based on this analysis, the Project Architect generates a comprehensive technical design that serves as the foundation for the entire pipeline.
  • Data Engineer: This agent handles the implementation of the data processing component, which involves tasks such as data standardization, augmentation, and transformation.
  • Model Engineer: This agent is tasked with selecting and configuring the model architecture by choosing the most appropriate base model, applying necessary modifications, and incorporating techniques such as regularization if needed.
  • Training Engineer: This agent is responsible for configuring the model training process, where key tasks include setting the hyperparameters, such as learning rate and batch size, and selecting the best optimization algorithms.
  • Performance Analyst: This agent is responsible for reviewing pipeline configurations and training logs after each training run is completed. It then identifies one area for improvement and provides targeted feedback to one selected engineer agent.
Each agent in the IMPROVE framework is equipped with knowledge of its role, the team’s goals, and how to collaborate effectively with other agents. Also, each agent operates with domain-specific expertise embedded through prompt engineering. For example, the Model Engineer is aware of effective pre-trained models like ConvNeXt (Liu et al., 2022), while the Data Engineer is familiar with advanced data augmentation techniques such as MixUp (Zhang et al., 2018) and CutMix (Yun et al., 2019). Although state-of-the-art LLMs implicitly understand these techniques and can implement them when asked to, they rarely apply them to code proactively unless explicitly instructed to do so. Thus, we specify these techniques in the prompts to ensure the agents only select and experiment with highly effective methods when generating the training pipeline.
The framework applies Iterative Refinement to optimize three primary pipeline components that we identified for object classification: data augmentation, model architecture, and training procedure. The development process begins with the Project Architect, who reviews the dataset and additional context to produce a technical design. This design is then passed to the Data Engineer, Model Engineer, and Training Engineer, who collaboratively build their respective components according to the plan. Once the data pipeline, model architecture, and training configurations are in place, the system integrates these components and trains the model. The Performance Analyst then reviews all past and current pipeline configurations, along with their associated training logs, to identify areas for improvement. It then provides targeted feedback to a selected engineer, who refines their component accordingly. The updated pipeline undergoes retraining and re-evaluation, ensuring that only changes leading to performance gains are retained, while ineffective modifications are logged for future reference but discarded from the active codebase. This iterative cycle continues for a predefined number of iterations, progressively refining the pipeline until the best-performing configuration and corresponding model checkpoint are identified as the final output.
3.4 Implementation
We now highlight some other design choices and innovations in IMPROVE that enhance its effectiveness beyond the Iterative Refinement strategy.
Dataset Information Utilization. A key advantage of the LLM agents framework is its ability to understand any natural language description of the dataset, enabling the system to make more informed decisions. Optionally, the user can supply a brief description of the dataset, detailing aspects such as the image domain (e.g., real-life objects, digits, synthetic images), whether the images are grayscale or colored, their dimensions (same or varying sizes), and their quality (high or low). This information is passed to both the Project Architect and Performance Analyst and is factored into the initial technical design and performance analysis. The impact of including this dataset information is discussed further in Section 4.3.
Autonomous Code Execution. The IMPROVE framework is capable of autonomously generating, executing, and refining model pipeline code over multiple iterations to improve model performance. To achieve this, IMPROVE first needs to ensure that the code produced by each engineer agent is coherent and compatible. This is facilitated by a predefined constant interface that specifies the required input arguments and expected return values for each component. The code-generating agents (i.e., Data Engineer, Model Engineer, and Training Engineer) have flexibility in their implementations but must adhere to these interface guidelines for input and output consistency. To automate the training process, a built-in workflow takes the agents’ responses, and then parses and extracts the code into Python files. The system utilizes the predefined interface to integrate these components into a coherent pipeline and initiates a terminal shell script that executes the full model training process. Training logs and other outputs are captured at the shell level and stored in a text file for analysis.
Unified Pipeline Initialization. At the start of the IMPROVE workflow, the Project Architect agent leverages the strong zero-shot generation capabilities of LLMs to generate a single, coherent pipeline in one step before dividing it into three components. In Section 4.3, we discuss how generating the data, model, and training components separately and then combining them instead can lead to suboptimal initial performance. Using unified pipeline initialization ensures greater efficiency by starting with a well-integrated baseline, reducing the need for extensive corrections and minimizing wasted iterations on poorly constructed initial configurations.
Summarized History. To ensure informed decision-making and prevent redundancy, all previous pipeline configurations and their corresponding training logs are passed to the Performance Analyst. However, passing the full history of code and logs could quickly exceed the LLM’s maximum context length. To mitigate this issue, after each iteration, the system makes a call to the LLM to generate a summary of the key aspects of the data, model, and training code, highlighting information such as the methods and hyperparameters used. Then, instead of passing the entire code from all past iterations, only the summarized configuration along with the training log is provided to the Performance Analyst to guide its reasoning.
@@ 4.1 Experimental Setting
We evaluated IMPROVE using two widely recognized classification datasets: CIFAR-10 (Krizhevsky and Hinton, 2009) and TinyImageNet (Deng et al., 2009). CIFAR-10 contains 60,000 32x32 color images across 10 distinct classes, while TinyImageNet, a subset of ImageNet1K, includes 100,000 images across 200 classes, resized to 64x64 pixels. Both datasets feature commonly seen objects such as vehicles and animals.
@@
  • Arabic Letters Classification (Khalil, 2023): This dataset contains 53,199 images of 65 written Arabic letters, each exhibiting positional variations based on their occurrence within a word with four possible forms: isolated, initial, medial and final. These letters were collected from 82 different users.
@@ 4.2 Main Results
4.3 Ablation Studies
@@
5 Conclusion
In this paper, we introduced Iterative Refinement, a novel strategy for designing LLM agent-based pipeline optimization systems by decomposing pipelines into independent components and refining each in isolation, enhancing interpretability and consistency while ensuring systematic improvements. To evaluate Iterative Refinement, we implemented IMPROVE, an end-to-end LLM agent framework that autonomously generates image classification models. Beyond Iterative Refinement, IMPROVE integrates several innovative features, including dataset-aware decision-making, automated code execution, and unified pipeline initialization, which together create a robust and adaptive optimization process. Our experiments demonstrated that IMPROVE consistently outperforms models generated through zero-shot LLM prompting, achieves results comparable to top human ML practitioners, and exhibits robustness against corrupted datasets while effectively adapting to diverse image domains. These findings highlight the potential of IMPROVE for real-world applications.

### v2

@@ 3.2 IMPROVE
To evaluate the Iterative Refinement strategy, we developed IMPROVE, a multi-agent end-to-end LLM framework that autonomously generates and optimizes an object classification pipeline.
We provide a diagram of the workflow of IMPROVE in Figure 1. It has five specialized roles.
  • Project Architect: Analyzes the dataset’s size, structure, and metadata to produce a comprehensive technical plan for the full pipeline.
  • Data Engineer: Implements data processing, including standardization, augmentation, and transformation routines.
  • Model Engineer: Selects and configures the model architecture, applying modifications and techniques such as regularization as needed.
  • Training Engineer: Sets hyperparameters (e.g., learning rate, batch size) and chooses appropriate optimization algorithms for model training.
  • Performance Analyst: Reviews pipeline configurations and training logs, then identifies one area for improvement and issues feedback to a selected engineer.
Each agent is aware of its responsibilities, the team’s overall goal, and how to collaborate effectively. Role-specific expertise is embedded through prompt engineering: for instance, the Model Engineer is aware of models like ConvNeXt (Liu et al., 2022), and the Data Engineer leverages advanced augmentations such as MixUp (Zhang et al., 2018) and CutMix (Yun et al., 2019). We explicitly included these techniques in the prompts because, while LLMs are capable of understanding and applying them when instructed, they rarely do so on their own.
@@
IMPROVE applies Iterative Refinement to optimize three primary components of image classification pipelines: data augmentation, model architecture, and training procedure. The workflow begins with the Project Architect generating a technical design from the dataset. This plan guides the Data Engineer, Model Engineer, and Training Engineer in constructing their respective modules. Once the pipeline is assembled, it is trained and evaluated. The Performance Analyst then reviews the results, identifies a single target for refinement, and prompts the corresponding engineer to revise their component. The system retrains and re-evaluates the updated pipeline, retaining only modifications that yield improved performance. Ineffective changes are discarded but logged for traceability. This loop continues for a fixed number of iterations, gradually converging to the best-performing pipeline and model checkpoint.
Engineering Specification
Beyond Iterative Refinement, IMPROVE integrates several engineering strategies for stable and autonomous pipeline construction.
Dataset-Aware Design. Users may optionally provide dataset metadata, such as domain, color mode, resolution, or quality, as a natural language description. This information is parsed by the Project Architect and Performance Analyst to inform initial design and diagnostics (see Section A.1).
Executable Agent Code. Engineer agents generate Python code conforming to predefined I/O interfaces, ensuring composability. IMPROVE parses and integrates these outputs into a unified script, executes the pipeline via shell, and logs results for downstream evaluation.
Unified Initialization. To improve early-stage stability, the Project Architect generates a complete pipeline in one zero-shot pass before delegating submodules. This yields stronger starting configurations than bottom-up assembly (see Section A.1).
Compact Iteration History. To avoid context overflow, each refinement step is summarized by an LLM. Only condensed configurations and logs are passed to the Performance Analyst, preserving key signals without exceeding input limits.
@@ 4.1 Experimental Setting
We evaluated IMPROVE using two widely recognized classification datasets: CIFAR-10 (Krizhevsky and Hinton, 2009) and TinyImageNet (Deng et al., 2009).
@@
  • Arabic Letters Classification (Khalil, 2023): This dataset contains 53,199 images of 65 written Arabic letters, each exhibiting positional variations based on their occurrence within a word with four possible forms: isolated, initial, medial and final, collected from 82 different users.
@@ 4.2 Main Results
5 Conclusion
We introduced Iterative Refinement, a strategy for LLM agent-based pipeline optimization that updates components independently for stable, interpretable improvements. We implemented this in IMPROVE, an end-to-end framework that automates image classification via modular agents. IMPROVE incorporates dataset-aware reasoning, code execution, and unified initialization. Experiments show that it outperforms zero-shot baselines, rivals expert performance, and adapts well to corrupted and diverse datasets, demonstrating its practical effectiveness.
Appendix A Ablation Studies
A.1 Ablation Studies
@@
Appendix B Proof
Proof of Theorem 3.1
Recall the cumulative improvement function:
$f(T):=\beta T-\frac{\alpha\gamma(1-\gamma^{T})}{1-\gamma}.$
We aim to show:
$\exists T^{*}\in\mathbb{N}\text{ such that }f(T^{*})\geq 0,\text{ and }f(T)\text{ is strictly increasing in }T.$
Step 1: Monotonicity.
We compute the discrete derivative:
$f(T+1)-f(T)=\beta-\alpha\cdot\gamma^{T+1}.$
Since $\gamma\in(0,1)$, we know $\gamma^{T+1}\to 0$ as $T\to\infty$. Therefore, for sufficiently large $T$, we have:
$f(T+1)-f(T)>0,$
implying that $f(T)$ is strictly increasing in $T$.
Step 2: Crossover point existence.
We analyze the behavior of $f(T)$ as $T\to\infty$:
$\lim_{T\to\infty}f(T)=\beta T-\frac{\alpha\gamma}{1-\gamma}.$
Since $\beta T$ grows linearly and the second term is constant, $f(T)\to\infty$. Moreover, $f(0)=-\frac{\alpha\gamma}{1-\gamma}<0$, so by monotonicity, there must exist a minimal $T^{*}\in\mathbb{N}$ such that $f(T^{*})\geq 0$.
$\blacksquare$
Proof of Lemma 3.2
We compare convergence rates under a fixed computational budget.
Step 1: Gradient vs. Coordinate Descent Rates.
Under strong convexity and smoothness assumptions:
GD: $f(x^{(T)})-f(x^{*})\leq\left(1-\frac{\mu}{L}\right)^{T}(f(x^{(0)})-f(x^{*})),$
IR (GS): $f(x^{(T)})-f(x^{*})\leq\left(1-\frac{\mu_{1}}{L_{\max}}\right)^{T}(f(x^{(0)})-f(x^{*})).$
Let $N_{\text{global}}$ and $N_{\text{local}}$ denote the number of steps under each strategy with equivalent cost.
Step 2: Cost Ratio.
Total token cost over $k$ iterations:
$g_{\text{global}}(k)$ $=c_{0}+pc_{1}+pc_{2}(k-1)=c_{0}+px,$
$g_{\text{local}}(k)$ $=c_{0}+c_{1}+c_{2}(k-1)=c_{0}+x,\quad\text{where }x:=c_{1}+c_{2}(k-1).$
Therefore:
$\frac{N_{\text{global}}}{N_{\text{local}}}=\frac{g_{\text{global}}}{g_{\text{local}}}=\frac{c_{0}+px}{c_{0}+x}.$
Step 3: Convergence Comparison.
We want:
$\left(1-\frac{\mu_{1}}{L_{\max}}\right)^{N_{\text{local}}}<\left(1-\frac{\mu}{L}\right)^{N_{\text{global}}}.$
Taking logarithms and rearranging:
$\frac{N_{\text{global}}}{N_{\text{local}}}<\frac{\log\left(1-\frac{\mu_{1}}{L_{\max}}\right)}{\log\left(1-\frac{\mu}{L}\right)}=\log_{1-\mu/L}\left(1-\frac{\mu_{1}}{L_{\max}}\right).$
Substituting the cost ratio:
$\frac{c_{0}+x}{c_{0}+px}<\log_{1-\mu/L}\left(1-\frac{\mu_{1}}{L_{\max}}\right).$
$\blacksquare$